### v5

@@ 4.2.3. Proposed improvements to AI safety benchmarks
The following suggestions are offered as constructive opportunities to complement existing benchmarks and to support broader governance use.
To address the capability gaps identified above, we suggest:
  1. (i)
    Data quality and transparency. Encourage diverse and well documented corpora across regions, languages, and demographics, with dataset cards or data statements that capture provenance, consent and licensing, and curation criteria to enhance traceability and accountability (Eriksson et al., 2025; Gebru et al., 2021).
  2. (ii)
    Construct clarity and validation. Make constructs explicit by stating what each task and metric is intended to measure and, where feasible, validate measures with multidisciplinary input to guard against superficial compliance claims (Ren et al., 2024).
  3. (iii)
    Evaluation resilience. Reduce opportunities for gaming or overfitting by incorporating dynamic or adversarial evaluation, including undisclosed items and evolving prompts, consistent with established security testing practices (Oprea and Vassilev, 2023; Ghosh et al., 2025).
  4. (iv)
    Broader scope. Expand coverage to multimodal, non-English, and higher-risk domains, and include multi-agent or autonomous settings where system level failures are more likely to surface (Salazar et al., 2025; Raza et al., 2025c).
  5. (v)
    Multi-stakeholder alignment. Mitigate potential commercial or domain bias by engaging multiple stakeholders and making explicit linkages to policy instruments and governance frameworks (Outeda, 2024).
  6. (vi)
    Timeliness and maintenance. Support continuous evaluation through automated pipelines that enable periodic re-testing across the model lifecycle, improving comparability over time (Wang et al., 2024).
  7. (vii)
    Exploration of unknown risks. Integrate structured red teaming and stress tests under distribution shift, and monitor for emergent behaviors that may elude static suites, drawing on existing practices such as ALERT and Rainbow Teaming (Tedeschi et al., 2024; Samvelyan et al., 2024; Liu et al., 2023a).
Figure 5. Responsible AI governance landscape and linkage to our evaluation criteria. Technical factors correspond primarily to rubric items C1–C7 (fairness/bias, toxicity, security/robustness, mis/disinformation, deepfakes/media integrity, privacy, system-level failure). Legal instruments (e.g., EU AI Act, NIST AI RMF) ground C10 (governance alignment). Risk & assurance processes (anticipation, responsiveness, reflexivity, inclusion, knowledge management) support adaptive testing (C8) and metric validity/record-keeping (C9). Sustainable AI highlights environmental/societal/economic impacts that are only weakly covered by current benchmarks, motivating the gaps observed in Tables 58.
@@
To operationalize these improvements and make evidence portable across audits, we now specify lifecycle KPIs that teams can track during development, deployment, and post-market monitoring.
4.2.4. Key Performance Indicators (KPIs) for Responsible AI
@@
KPIs provide measurable criteria to assess the performance of AI systems across critical dimensions of responsibility (Mattioli et al., 2024). Unlike ethical principles that are mostly subjective, KPIs translate RAI goals into operational metrics that can be tracked, compared, and audited across the model lifecycle. They bridge governance expectations with technical implementation so that fairness, transparency, robustness, privacy, and sustainability are evidenced through quantifiable measures, enabling compliance and continuous improvement. We adapt established measures (SPD, DI, accuracy parity, robustness under perturbation, and energy) from prior work (Pessach and Shmueli, 2022; Rajkomar et al., 2018; Van Wynsberghe, 2021), and introduce pragmatic, author-defined KPIs for explainability, inclusivity, auditing cadence, and high-stakes errors to support governance-ready evidence.
@@
Notation and conventions. $f(x)$ is the model score for input $x$ (state whether probability, logit, or other monotone score); $y\in\{0,1\}$ is the true label; $\hat{y}=\mathbb{1}\{f(x)\geq t\}$ is the predicted label with a single threshold $t$ applied to all subgroups unless stated; $A$ is a protected attribute with subgroups $a$ and $b$; $\phi_{0}$ is a baseline output and $\phi_{i}(x)$ is the attribution for feature $i$ (e.g., SHAP or Integrated Gradients); “accuracy” refers to a stated metric (e.g., top-1, F1, AUROC) computed per subgroup with the same threshold $t$; a perturbation is a noise or attack consistent with a stated threat model; $P_{\mathrm{avg}}$ is average power (W), $T$ is time (h), $\mathrm{CI}$ is grid carbon intensity (kgCO2e/kWh), and $\mathrm{PUE}$ is Power Usage Effectiveness.
@@
    It defines the difference in positive outcome rates across subgroups. We specify state $A$, $t$, sample sizes; or bootstrap and a significance test.
@@
    This measures the absolute gap in a stated accuracy metric across subgroups (Raza et al., 2024c). It reports per subgroup and key intersections; keep $t$ fixed across subgroups.
@@
    It measures the error rate and accuracy drop under stated threat models (Raza et al., 2025d). We specify attack type and strength (e.g., $\ell_{p}$ bounds), number of steps, and per-attack results.
@@
    This measures incidence of errors that trigger a predefined harm tier (Kunreuther et al., 2002). We define the harm taxonomy and decision costs; consider a harm weighted rate in safety critical settings.
@@ 4.3. Explainable AI
Having established how benchmarks cover risk families, we examine explainability (XAI) methods in GenAI methods and what they can and can not do, evidence under audit. Explainable AI (XAI) comprises methods that render model behaviour and outputs intelligible to human stakeholders (Meske et al., 2022). Effective explanations serve multiple roles: they support accountability and trust, enable bias and safety analyses, and facilitate compliance with emerging governance requirements (Baker and Xiang, 2023; Arrieta et al., 2020). In practice, XAI must operate across technical, social, legal, and organisational layers rather than as a purely algorithmic add-on.
Figure 6. Post-hoc explainability within a Responsible AI (RAI) pipeline. A black-box model produces predictions $\hat{y}$ and attribution signals $I(x_{i})$, where $I(x_{i})$ denotes a feature-importance signal (e.g., gradient/saliency or a SHAP value). Interpretability pathways (feature relevance, example-based, surrogates, and visual/text explanations) feed fairness diagnostics (e.g., SPD/DI) and calibration analyses, supporting governance-aligned optimisation $\mathcal{L}_{\mathrm{RAI}}=\mathcal{L}+\lambda\,\mathcal{L}_{\mathrm{fair}}$. For LLMs/VLMs, prompt traces and cross-modal attention provide additional observability.
Scope and challenges in the GenAI era.
@@
Methods in practice.
XAI methods can be ante-hoc (interpretability built into the model class, e.g., sparse linear models, rule lists, decision trees) or post-hoc (explain an existing black-box locally or globally). Common post-hoc families include feature relevance (saliency/attribution, gradients, SHAP/LIME) for instance-level insight (Lundberg and Lee, 2017; Ribeiro et al., 2016a); example-based explanations (prototypes, nearest neighbours) that ground behaviour in the data manifold (Kenny et al., 2021); surrogate models that approximate decision logic with a transparent model; and visual/textual rationales (heatmaps, attention summaries) for human inspection (He et al., 2023a). For LLMs and Vision-language-models (VLMs), these are complemented by prompt-engineering probes, chain-of-thought (CoT) audits, and tool-use traces (Natarajan and Nambiar, 2024). Importantly, attention maps and saliency are not proofs of causal reasoning and should be triangulated with counterfactual tests and ablations (Jain and Wallace, 2019; Serrano and Smith, 2019).
Figure 6 situates post-hoc explanations within an RAI pipeline. Predictions $\hat{y}=M_{\phi}(x)$ can be accompanied by gradient-based importance $I(x_{i})=\partial M_{\phi}/\partial x_{i}$ to support local inspection; example-based evidence provides case-level comparators; and visual/text explanations aid end-user understanding. Fairness is evaluated with standard disparity measures such as Statistical Parity Difference (SPD) (Kun, 2015) and Disparate Impact (DI) (Barocas and Selbst, 2016):
@@
A compact theoretical frame.
Let $f_{\theta}:\mathcal{X}\!\to\!\mathcal{Y}$ denote a predictive model (or $p_{\theta}(\cdot\mid x,c)$ for generative systems with context $c$). We define an explanation operator
@@
Desiderata (axioms). An explanation $z=\mathcal{E}(x,f_{\theta},\omega)$ is useful when it approximately satisfies: (i) local accuracy (additive decompositions sum to output deviation from a baseline) (Lundberg and Lee, 2017); (ii) sensitivity/monotonicity (if $f_{\theta}$’s dependence on feature $i$ increases at $x$, its attribution should not decrease) (Sundararajan et al., 2017); (iii) invariance to representation-preserving transforms and to behaviourally equivalent models (Sundararajan et al., 2017); (iv) stability (Lipschitz continuity in $x$ and $\omega$(Alvarez-Melis and Jaakkola, 2018); (v) counterfactual consistency (factors highlighted by $z$ admit feasible interventions that flip or materially shift the output); and (vi) parsimony (minimal complexity $\Omega(z)$ subject to fidelity) (Kusner et al., 2017). These properties specialise naturally to multimodal settings via alignment between rationales and grounded visual evidence.
@@
Limitation: Even theoretically sound explanations can mislead when they are plausible but unfaithful (e.g., over-reading attention maps), trigger Goodhart effects (Strathern, 1997b) (optimising the appearance of fairness rather than outcomes), or lack distributional validity (narratives drift across domains). Reporting stability, counterfactual checks, and provenance mitigates these risks and makes explanations suitable as audit evidence rather than marketing artefacts.
Next, we present AI-ready test-beds that computes lifecycle KPIs and logging XAI evidence as audit-ready artifacts.
@@ 4.4. AI-ready Test-beds
AI-ready test-beds advance RAI by providing controlled environments to develop, test, and validate AI models with reliability and reproducibility guarantees. In the post-ChatGPT era, where GenAI is increasingly embedded in products and services, such platforms help assess non-determinism, multimodal behavior, and risks like bias, hallucination, and privacy leakage, supporting proportionate assurance and governance alignment. We present representative examples in Table LABEL:tab:ai_testbeds. The key value of these test-beds is in producing audit-ready evidence datasheets/model cards, prompt/tool logs, KPI dashboards, red-team logs, and versioned runs.
Table 9. AI evaluation platforms and initiatives used as de facto test-beds for RAI. We distinguish formal test-beds/platforms from broader labs/initiatives commonly used for evaluation.
@@
AI4EU AI-on-Demand Platform (cor, 2019) General AI development Platform Ethical-by-design resources; model/dataset catalogue; support for transparency and explainability workflows (used to prototype responsible AI pipelines).
IEEE Ethical AI Systems Test Bed (win, 2019) Ethical AI development Test-bed/Platform Evaluation against ethical frameworks; human-centred design; fairness-oriented assessment protocols; documentation of decision rationale.
AI Testbed for Trustworthy AI (TNO) (tno, 2020) Trustworthy AI Test-bed Robustness and transparency checks; fairness analysis; risk/assurance reporting for deployment contexts.
ETH Zurich Safe AI Lab (saf, 2020) Safe and fair AI Lab/Platform Safety-critical validation; robustness/hallucination probing; reliability and interpretability methods for high-stakes settings.
HUMANE AI (hum, 2020) Human-centric AI Platform/Initiative Alignment with human values; societal-impact and fairness guardrails; guidance for participatory, responsible design.
AI for Good (ITU) (AIf, 2017) Social good / RAI Initiative/Program UN SDG–aligned use cases; ethics and societal-benefit framing; showcases reproducible, responsible deployments.
UKRI Trustworthy Autonomous Systems (TAS) Hub (TAS, 2020) Autonomous systems Program/Test-bed Accountability and transparency for autonomy; safety cases; conformance/compliance support for real-world systems.
@@
ClarityNLP (Healthcare) (Cla, 2021) Healthcare NLP Platform Clinical text pipelines; emphasis on fairness, transparency, and ethical data use; supports domain-specific evaluation.
@@
Chatbot Arena (LMSYS) (cha, 2024) LLM evaluation Community platform Open-source platform where users compare LLMs in head-to-head “arena” battles; pairwise ranking with more than 800k community votes and transparent leaderboards; all evaluation pipelines and data are public for reproducibility.
AI Fairness 360 (AIF360) (Bellamy et al., 2018a) Fairness evaluation Toolkit Open-source Python toolkit with 70+ fairness metrics and 9 mitigation algorithms; supports bias detection, fairness explanations and interactive web interface for non-experts.
Zeno (zen, 2023) AI evaluation Platform Interactive evaluation platform; enables data exploration across any data type and automatically discovers systematic failures using slice finder; supports interactive charts to compare evaluation metrics across slices and models.
Galileo (gal, 2023) Generative AI evaluation Platform Automated evaluation metrics for GenAI (RAG and agent metrics plus safety/security metrics); real-time guardrails to block hallucinations, PII and prompt injection; captures corner cases and allows custom metrics.
@@
PromptLayer (pro, 2022) Prompt engineering Platform Prompt-evaluation platform with visual pipelines for scoring prompts against golden datasets, backtesting with historical data, regression testing and continuous integration of evaluation pipelines; useful for systematic RAG/agent evaluations.
@@
5. RAI Application Across Domains
GenAI systems are increasingly embedded in decision-making processes across diverse domains such as healthcare, finance, human resource, defense and education. While these systems offer significant utility, they also introduce new categories of risks and challenges, as outlined earlier. The challenges and mitigation strategies synthesized in this review are broadly relevant, but their salience and priority vary substantially by sectoral context. We translate the above findings into stakeholder-aligned practices and a deployment-oriented roadmap, with examples and decision pointers.

### v6

@@ 4.2.3. Proposed Improvements to AI Safety Benchmarks
Based on the benchmark gaps we found in the literature, we suggest constructive opportunities to complement existing benchmarks and to support broader governance use. First,we recommend that benchmark developers encourage diverse and well-documented corpora across regions, languages, and demographics, accompanied by dataset cards or data statements that capture provenance, consent, licensing, and curation criteria to enhance traceability and accountability (Eriksson et al., 2025; Gebru et al., 2021). Equally important is construct clarity: each task and metric should state explicitly what it is intended to measure and, where feasible, be validated with multidisciplinary input to guard against superficial compliance claims (Ren et al., 2024). Without these foundations, high benchmark scores risk measuring test-taking ability rather than genuine safety.
Incorporating adversarial evaluation, including undisclosed items and evolving prompts can be used bring benchmarks in line with established security testing practices (Oprea and Vassilev, 2023b; Ghosh et al., 2025). Scope should also expand to cover multimodal, non-English, and higher-risk domains, as well as multi-agent or autonomous settings where system-level failures are more likely to surface (Salazar et al., 2025). These benchmarks rarely make explicit linkages to policy instruments. Engaging multiple stakeholders and mapping evaluation outputs to governance frameworks would mitigate commercial or domain bias and improve regulatory utility (Outeda, 2024). Continuous evaluation through automated pipelines that enable periodic re-testing across the model lifecycle would further improve comparability over time (Wang et al., 2024). Finally, structured red teaming and stress tests under distribution shift should be integrated to surface emergent behaviours that may elude static suites, drawing on practices such as ALERT and Rainbow Teaming.
A broader concern emerging from this analysis is that many existing benchmarks may facilitate what might be termed “safetywashing”: models can achieve high scores on structured benchmarks such as SafetyBench (Zhang et al., 2024) by learning keyword-driven refusal scripts, while failing under the multi-turn, emotionally complex interactions documented in real-world deployment incidents (Roose, 2023). This pattern underscores the need to move beyond static compliance checks toward evaluation paradigms that capture the adversarial and socially situated nature of human–model interaction.
@@
The governance landscape framework is intended to illustrate that responsibility is not a single algorithmic check, but a multi-dimensional stewardship involving technical, legal, and social factors. Figure 8 provides the foundational logic for our ten-criterion rubric, ensuring that our technical assessment is grounded in real-world governance requirements. To operationalize these improvements and make evidence portable across audits, we now specify lifecycle KPIs that teams can track during development, deployment, and post-market monitoring.
4.2.4. Key Performance Indicators (KPIs) for Responsible GenAI
@@
KPIs provide measurable criteria to assess the performance of AI systems across critical dimensions of responsibility (Mattioli et al., 2024). Unlike ethical principles that are mostly subjective, KPIs translate RAI goals into operational metrics that can be tracked, compared, and audited across the model lifecycle. They bridge governance expectations with technical implementation so that fairness, transparency, robustness, privacy, and sustainability are evidenced through quantifiable measures, enabling compliance and continuous improvement. We adapt established measures (SPD, DI, accuracy parity, robustness under perturbation, and energy) from prior work (Pessach and Shmueli, 2022), and introduce pragmatic, author-defined KPIs for explainability, inclusivity, auditing cadence, and high-stakes errors to support governance-ready evidence.
@@
Notation and conventions $f(x)$ is the model score for input $x$ (state whether probability, logit, or other monotone score); $y\in\{0,1\}$ is the true label; $\hat{y}=\mathbb{1}\{f(x)\geq t\}$ is the predicted label with a single threshold $t$ applied to all subgroups unless stated; $A$ is a protected attribute with subgroups $a$ and $b$; $\phi_{0}$ is a baseline output and $\phi_{i}(x)$ is the attribution for feature $i$ (e.g., SHAP or Integrated Gradients); “accuracy” refers to a stated metric (e.g., top-1, F1, AUROC) computed per subgroup with the same threshold $t$; a perturbation is a noise or attack consistent with a stated threat model; $P_{\mathrm{avg}}$ is average power (W), $T$ is time (h), $\mathrm{CI}$ is grid carbon intensity (kgCO2e/kWh), and $\mathrm{PUE}$ is Power Usage Effectiveness.
@@
[…]
@@
    This measures the absolute gap in a stated accuracy metric across subgroups (Raza et al., 2024). It reports per subgroup and key intersections; keep $t$ fixed across subgroups.
@@
    It measures the error rate and accuracy drop under stated threat models. We specify attack type and strength (e.g., $\ell_{p}$ bounds), number of steps, and per-attack results. For generative models, robustness is more appropriately measured as attack success rate (ASR) under jailbreak or prompt-injection attempts, safety-refusal consistency under paraphrase, and output stability under temperature and seed variation, rather than misclassification rate under $\ell_{p}$ perturbations
@@
    This measures incidence of errors that trigger a predefined harm tier (Kunreuther et al., 2002). We define the harm taxonomy and decision costs; consider a harm weighted rate in safety critical settings. In generative deployments, critical errors include hallucinated medical or legal facts, harmful completions, and PII leakage incidents; $E_{h}$ should accordingly be computed as the incidence of such events per $N$ generations in the target domain, with harm severity tiers defined by deployment context
@@
Sections 56 use this rubric (C1–C10) and the lifecycle KPIs as a running lens to interpret domain evidence and highlight where current practice still lacks governance-ready coverage.
@@ 4.3. Explainable AI
Recent GenAI era deployments expose failure modes (hallucination, privacy leakage, deepfakes, and tool-use errors) that cut across domains; we therefore use C1–C10 and the KPI set as a consistent lens throughout the remainder of the survey. Having established how benchmarks cover risk families, we examine explainability methods and what they can and can not do, evidence under audit. Explainable AI (XAI) comprises methods that render model behaviour and outputs intelligible to human stakeholders (Meske et al., 2022). Effective explanations serve multiple roles: they support accountability and trust, enable bias and safety analyses, and facilitate compliance with emerging governance requirements (Baker and Xiang, 2023; Arrieta et al., 2020). In practice, XAI must operate across technical, social, legal, and organisational layers rather than as a purely algorithmic add-on.
Figure 9. Post-hoc explainability within a Responsible AI (RAI) pipeline. A black-box model produces predictions $\hat{y}$ and attribution signals $I(x_{i})$, where $I(x_{i})$ denotes a feature-importance signal (e.g., gradient/saliency or a SHAP value). Interpretability pathways (feature relevance, example-based, surrogates, and visual/text explanations) feed fairness diagnostics (e.g., SPD/DI) and calibration analyses, supporting governance-aligned optimisation $\mathcal{L}_{\mathrm{RAI}}=\mathcal{L}+\lambda\,\mathcal{L}_{\mathrm{fair}}$. For LLMs/VLMs, prompt traces and cross-modal attention provide additional observability
Scope and challenges in the GenAI
@@
Methods in practice
XAI methods can be ante-hoc (interpretability built into the model class, e.g., sparse linear models, rule lists, decision trees) or post-hoc (explain an existing black-box locally or globally). Common post-hoc families include feature relevance (saliency/attribution, gradients, SHAP/LIME) for instance-level insight (Lundberg and Lee, 2017; Ribeiro et al., 2016a); example-based explanations (prototypes, nearest neighbours) that ground behaviour in the data manifold (Kenny et al., 2021); surrogate models that approximate decision logic with a transparent model; and visual/textual rationales (heatmaps, attention summaries) for human inspection (He et al., 2023). For LLMs and Vision-language-models (VLMs), these are complemented by prompt-engineering probes, chain-of-thought (CoT) audits, and tool-use traces (Natarajan and Nambiar, 2024). Importantly, attention maps and saliency are not proofs of causal reasoning and should be triangulated with counterfactual tests and ablations (Jain and Wallace, 2019).
Figure 9 situates post-hoc explanations within an RAI pipeline. This framework bridges the gap between abstract interpretability theory and concrete engineering practice. Its specific role is to show how attribution signals and fairness checks (Barocas and Selbst, 2016) can be integrated into a standard pipeline to produce the ”portable evidence” needed for external audits and institutional trust. Predictions $\hat{y}=M_{\phi}(x)$ can be accompanied by gradient-based importance $I(x_{i})=\partial M_{\phi}/\partial x_{i}$ to support local inspection; example-based evidence provides case-level comparators; and visual/text explanations aid end-user understanding. Fairness is evaluated with standard disparity measures such as Statistical Parity Difference (SPD) (Kun, 2015) and Disparate Impact (DI) (Barocas and Selbst, 2016):
@@
A compact theoretical frame
Let $p_{\theta}(\cdot\mid x,c,\omega)$ denote a generative model with input $x$, context $c$ (e.g., system prompt, retrieved documents), and stochastic decoding parameters $\omega$ (temperature, top-$k/p$). The predictive case $f_{\theta}:\mathcal{X}\to\mathcal{Y}$ is recovered as a deterministic special instance
@@
Desiderata (axioms) An explanation $z=\mathcal{E}(x,f_{\theta},\omega)$ is useful when it approximately satisfies: (i) local accuracy (additive decompositions sum to output deviation from a baseline) (Lundberg and Lee, 2017); (ii) sensitivity/monotonicity (if $f_{\theta}$’s dependence on feature $i$ increases at $x$, its attribution should not decrease) (Sundararajan et al., 2017); (iii) invariance to representation-preserving transforms and to behaviourally equivalent models (Sundararajan et al., 2017); (iv) stability (Lipschitz continuity in $x$ and $\omega$(Alvarez-Melis and Jaakkola, 2018); (v) counterfactual consistency (factors highlighted by $z$ admit feasible interventions that flip or materially shift the output); and (vi) parsimony (minimal complexity $\Omega(z)$ subject to fidelity) (Kusner et al., 2017). These properties specialise naturally to multimodal settings via alignment between rationales and grounded visual evidence.
@@
Limitation: Even theoretically sound explanations can mislead when they are plausible but unfaithful (e.g., over-reading attention maps), trigger Goodhart effects (Strathern, 1997b) (optimising the appearance of fairness rather than outcomes), or lack distributional validity (narratives drift across domains). Reporting stability, counterfactual checks, and provenance mitigates these risks and makes explanations suitable as audit evidence rather than marketing artefacts. Based on our synthesis findings, there is a dangerous ”Explanation Gap” in how we use these tools. While SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016b) provide useful snapshots, the literature reveals they often create ”plausible but unfaithful” stories. They give us a narrative that makes sense to a human, but because transformers are so complex and non-linear, these stories often don’t match what is actually happening under the hood (Strathern, 1997a). This creates a false sense of security that we think we’ve reached transparency when we’ve really just seen a surface-level summary. Next, we present AI-ready test-beds that computes lifecycle KPIs and logging XAI evidence as audit-ready artifacts.
@@ 4.4. AI-ready Test-beds
AI-ready test-beds advance RAI by providing controlled environments to develop, test, and validate AI models with reliability and reproducibility guarantees. In the post-ChatGPT era, such platforms help assess non-determinism, multimodal behavior, and risks like bias, hallucination, and privacy leakage, supporting proportionate assurance and governance alignment. We present representative examples in Table LABEL:tab:ai_testbeds. The key value of these test-beds is in producing audit-ready evidence datasheets/model cards, prompt/tool logs, KPI dashboards, red-team logs, and versioned runs.
Table 7. AI evaluation platforms and initiatives used as de facto test-beds for RAI. We distinguish formal test-beds/platforms from broader labs/initiatives commonly used for evaluation
@@
AI4EU AI-on-Demand Platform (AI4EU, 2019) General AI development Platform Ethical-by-design resources; model/dataset catalogue; support for transparency and explainability workflows (used to prototype responsible AI pipelines).
IEEE Ethical AI Systems Test Bed (IEEE, 2019) Ethical AI development Test-bed/Platform Evaluation against ethical frameworks; human-centred design; fairness-oriented assessment protocols; documentation of decision rationale.
AI Testbed for Trustworthy AI (TNO) (for Applied Scientific Research, 2020) Trustworthy AI Test-bed Robustness and transparency checks; fairness analysis; risk/assurance reporting for deployment contexts.
ETH Zurich Safe AI Lab (Zurich, 2020) Safe and fair AI Lab/Platform Safety-critical validation; robustness/hallucination probing; reliability and interpretability methods for high-stakes settings.
HUMANE AI (HumaneAI, 2020) Human-centric AI Platform/Initiative Alignment with human values; societal-impact and fairness guardrails; guidance for participatory, responsible design.
AI for Good (ITU) ((2017), ITU) Social good / RAI Initiative/Program UN SDG–aligned use cases; ethics and societal-benefit framing; showcases reproducible, responsible deployments.
UKRI Trustworthy Autonomous Systems (TAS) Hub (Research and Innovation, 2020) Autonomous systems Program/Test-bed Accountability and transparency for autonomy; safety cases; conformance/compliance support for real-world systems.
@@
ClarityNLP (Healthcare) (ClarityNLP, 2021) Healthcare NLP Platform Clinical text pipelines; emphasis on fairness, transparency, and ethical data use; supports domain-specific evaluation.
@@
Chatbot Arena (LMSYS) (Org, 2024) LLM evaluation Community platform Open-source platform where users compare LLMs in head-to-head “arena” battles; pairwise ranking with more than 800k community votes and transparent leaderboards; all evaluation pipelines and data are public for reproducibility.
AI Fairness 360 (AIF360) (Bellamy et al., 2018) Fairness evaluation Toolkit Open-source Python toolkit with 70+ fairness metrics and 9 mitigation algorithms; supports bias detection, fairness explanations and interactive web interface for non-experts.
Zeno (Zeno, 2023) AI evaluation Platform Interactive evaluation platform; enables data exploration across any data type and automatically discovers systematic failures using slice finder; supports interactive charts to compare evaluation metrics across slices and models.
Galileo (Labs, 2023) Generative AI evaluation Platform Automated evaluation metrics for GenAI (RAG and agent metrics plus safety/security metrics); real-time guardrails to block hallucinations, PII and prompt injection; captures corner cases and allows custom metrics.
@@
PromptLayer (PromptLayer, 2022) Prompt engineering Platform Prompt-evaluation platform with visual pipelines for scoring prompts against golden datasets, backtesting with historical data, regression testing and continuous integration of evaluation pipelines; useful for systematic RAG/agent evaluations.
@@
4.4.1. The Lifecycle Evidence Matrix: Mapping Metrics to Engineering Controls
To facilitate the implementation of our proposed evaluation infrastructure, we synthesize the relationship between KPIs, benchmarks, and technical tools. As argued throughout this survey, the effectiveness of Responsible GenAI hinges on the ability to produce portable, verifiable evidence across the lifecycle. Table 8 provides a decision-support mapping that links the KPIs defined in Section 4.2.4 to the specific testbeds and datasets that operationalize them. This matrix enables practitioners to select the appropriate ”tooling stack” based on their specific governance requirements, whether prioritizing fairness audits in Finance or robustness testing in Defense.
Table 8. The ResGenAI Accountability Matrix: Linking KPIs to Benchmarks, Datasets, and Audit Toolkits
KPI Category Primary KPIs Benchmarks & Datasets Audit Toolkits & Testbeds
Fairness & Bias SPD, DI, $\Delta_{\mathrm{acc}}$ SafetyBench (Zhang et al., 2024), HumaniBench (Raza et al., 2025) AIF360 (Bellamy et al., 2018), Fairlearn (Weerts et al., 2023), Arize (Dong et al., 2024)
Safety & Toxicity $E_{h}$ (High-stakes error) HarmBench (Mazeika et al., 2024), SALAD-Bench (Li et al., 2024a) Unitary (Ltd., [n. d.]), Project Moonshot (AI Verify Foundation, 2024), Galileo (Labs, 2023)
Robustness & Security $R$ (Robustness), ASR ALERT (Tedeschi et al., 2024), Rainbow Teaming (Samvelyan et al., 2024) ART (Research, [n. d.]), PromptLayer (PromptLayer, 2022), ToolSandbox (Lu et al., 2024a)
Privacy $P$ (Compliance rate) PrivLM-Bench (Li et al., 2024b), PrivacyLens (Shao and et al., 2024) Privacy Meter (Privacy and Lab, [n. d.]), Google DP (Google, [n. d.]), SecretFlow (Group, [n. d.])
Explainability $E(x)$, LFG (Faithfulness) HELM (Liang et al., 2023), XSTest (Röttger et al., 2024) LIT (Google PAIR, [n. d.]), InterpretML, AIX360 (Arya et al., 2019)
Sustainability $E_{\mathrm{kWh}}$, $C_{\mathrm{CO_{2}e}}$ CodeCarbon (Song et al., 2025) Evidently (Evidently AI, [n. d.]), Arize (Dong et al., 2024), Zeno (Zeno, 2023)
4.4.2. Turning Metrics into Assurance: The ResGenAI Audit Loop
The technical efficacy of ResGenAI remains localized unless it can be translated into portable audit evidence. As visualized in Fig. 10, our framework proposes a three layer loop that bridges the Evaluation Gap identified in Section 6.4. In this cycle, the technical KPIs defined in Section 4.2.4 act as the raw sensory data for the ecosystem. These metrics are packaged into standardized audit artifacts such as model cards and red teaming logs which serve as the portable evidence required by compliance auditors under the EU AI Act (European Union, 2024) or NIST AI RMF (of Standards and, NIST). By establishing this formal “chain of trust”, organizations can move beyond static safety checks toward continuous stewardship where regulatory feedback directly informs the recalibration of technical metrics.
Figure 10. The ResGenAI Audit Loop: translating technical KPIs into portable audit evidence to support regulatory assurance and continuous stewardship across the foundation model lifecycle.
5. Responsible Generative AI Application Across Domains
GenAI systems are increasingly embedded in decision-making processes across diverse domains such as healthcare, finance, human resource, defense and education. While these systems offer significant utility, they also introduce new categories of risks and challenges, as outlined earlier. The challenges and mitigation strategies synthesized in this review are broadly relevant, but their salience and priority vary substantially by sectoral context. We translate the above findings into stakeholder-aligned practices and a deployment-oriented roadmap, with examples and decision pointers. Each subsection below examines domain-specific applications, challenges, and strategies through the lens of our governance-aligned rubric (C1–C10, Section 4.2.1) and lifecycle KPIs (Section 4.2.4). For every domain we (i) identify which rubric criteria are most salient, (ii) assess how the benchmarks in Tables 46 cover those risks, and (iii) indicate which KPIs practitioners should prioritize. Table 9 consolidates this analysis, mapping each sector to its primary criteria gaps and priority KPIs.